OpenStudy (anonymous):

Solve the equation 15(x - 3) = 12x - 43. A. 2/3 B. 3/2 C. -40/3 D. 3

OpenStudy (anonymous):

first distribute 15 on the left side 15x-45=12x-43 can you now solve this ? just bring the x terms to one side and constant terms to ther sides.

OpenStudy (anonymous):

\[ 15(x - 3) = 12x - 43 \\ 15x -45 = 12x - 43 \]

OpenStudy (anonymous):

\[15x−45=12x−43 \\ 15x−45+45=12x−43+45 \\ 15x=12x+2 \\ 15x-12x=12x-12x+2 \\ 3x=2\]

OpenStudy (anonymous):

divide both sides by 3 and you have your answer